If the area A of a triangle is 60 sq m and the base b is 20 m what is the altitude h?
Answer by nerdybill(7384)

If the area A of a triangle is 60 sq m and the base b is 20 m what is the altitude h?
.
Area of any triangle is given by the formula:
A = (1/2)bh
where
A is the area
b is the base
h is the height or altitude
.
Plug in what was given and solve for h:
A = (1/2)bh
60 = (1/2)(20)h
60 = (10)h
6 meters = h